OpenStudy (anonymous):

You and your sister decide to combine your weekly overtime earnings to buy a birthday gift for another mother. Your overtime rate is $18 per hour and your sisters overtime rate is $24 per hour. The total amount earned for the gift was$288. If you worked two more hours of overtime than our sister. How many overtime hours did each of you work?

OpenStudy (ikram002p):

so if ur sis over time hour = x then ur over time hour =x+2 18 (x+2)+ 24 x= 288 18x+36+24x=288 42 x=252 x=6 so total hour is x+(x+2)=6+8=14

OpenStudy (ikram002p):

got it nw ?